OpenStudy (anonymous):

can somebody please show me how to solve this convert the angle measure to decimal degrees, the nearest hundredth of a degree. 25° 25' 40''

OpenStudy (anonymous):

You have 25 minutes and 40 seconds (in addition to the 25 degrees). 25 minutes x (1 degree / 60 minutes) + 40 seconds x (1 degree / 3600 seconds) + 25 degrees

OpenStudy (mathstudent55):

Use the same conversion factors you had in the previous problems: 1 degree = 60 minutes 1 minute = 60 seconds Start with the seconds amount: 40 seconds * (1 minute/60 seconds) * (1 degree/60 minutes) = 40/(60 * 60) degrees 25 minutes * (1 degree/60 minutes) = 25/60 degrees 25 degrees = 25 degrees Now do the operations and add all the right sides together.

OpenStudy (anonymous):

So, you have 25 + 25/60 + 40/3600

OpenStudy (anonymous):

This simplifies to: 25 + 5/12 + 1/90

OpenStudy (anonymous):

thank you both so much

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @erin512 And you're welcome!

OpenStudy (anonymous):

what would the final answer be in this case @tcarroll010? 25 degrees and what else?

OpenStudy (anonymous):

np, continuing on: 25 + 75/180 + 2/180 25 + 77/180 25 + 0.4277777777777 25.43

OpenStudy (anonymous):

25.43 is the answer to what?

OpenStudy (anonymous):

To the number of decimal degrees to the nearest hundredth. Unless there is something more to the problem that is not stated in the problem.

OpenStudy (anonymous):

oh ok, so is that the final answer or is there more than 1? :)

OpenStudy (anonymous):

That should be it. Do you have an answer key to check it against?

OpenStudy (anonymous):

no but this is a practice assignment and once i submit it i can get the answer key to check it with :) thanks so much for the explanation!!!

OpenStudy (anonymous):

you're quite welcome and have a nice day!

OpenStudy (anonymous):

@tcarroll010 if i post one more of these can you help explain it again just so i can get a better understanding? its good practice for me, thanks so much :)

OpenStudy (anonymous):

ill post it now

OpenStudy (anonymous):

ok